OpenStudy (anonymous):

Partial Fraction indef. integral of: (e^x)/(1-e^(3x)) tried wolframalpha, and got to the u substitution.. it got really complicated after that.. and I need help. Thanks!

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ e^x }{ 1-e ^{3x} }\]

OpenStudy (anonymous):

forgot the dx at the end of the integral, sorry

myininaya (myininaya):

How would you use partial fractions to decompose this fraction: \[\frac{u}{1-u^3}\]

OpenStudy (anonymous):

change the sign to negative.. giving you negative integral of u/(u^3-1)

myininaya (myininaya):

you should be able to factor a difference of cubes

OpenStudy (anonymous):

making it a polynomial.. which i got.. but after solving and finding out what a, bx and c are, you have to repeat?

OpenStudy (anonymous):

oh, yeah.. that's (u-1)(u^2 + u + 1)

myininaya (myininaya):

yep so one of those factors is linear and the other factor is an irreducible quadratic we are done with the factoring bit

myininaya (myininaya):

for linear factors, you do constant/linear for irreducible quadratics, you do (constant*the variable +another constant)/the quadratic

myininaya (myininaya):

\[\frac{-u}{u^3-1}=\frac{A}{u-1}+\frac{Bu+C}{u^2+u+1}\]

OpenStudy (anonymous):

ok, for A i got 1/3 (Bu+ C )is ((1/3)u + 2/3)

OpenStudy (anonymous):

integrating the first half i get (1/3)ln|u| = (1/3)*ln|e^x|

myininaya (myininaya):

I'm not getting that for C... Let me recheck.

myininaya (myininaya):

oh and I got A=-1/3 but let me recheck

OpenStudy (anonymous):

oh, i took the negative out.. if that changes anything/

myininaya (myininaya):

So you are going to multiply by the negative later.

OpenStudy (anonymous):

yeah.. ok so A = 1/3 B = -1/3 C= -2/3

OpenStudy (anonymous):

yeah.. it's less confusing for me to take the negative out and multiply by it at the end

myininaya (myininaya):

So you did this then? \[u=A(u^2+u+1)+(Bu+C)(u-1)\] \[u=A(u^2+u+1)+(Bu^2-Bu+Cu-C)\] \[u=(A+B)u^2+(A-B+C)u+(A-C)\]

myininaya (myininaya):

\[=>A+B=0 \text{ and } A-B+C=1 \text{ and } A-C=0\]

myininaya (myininaya):

I'm just decomposing the fraction. We can come back to the integral later.

myininaya (myininaya):

I want to see what the integral looks like without substitution at first because she said she didn't want substitution.

myininaya (myininaya):

We can make substitution later if we need to.

OpenStudy (anonymous):

yeah, when solving for A/B/C don't you set it equal to the numerator? so if it were u.. that would mess things up

myininaya (myininaya):

We are looking at the fraction part right now.

myininaya (myininaya):

ok anyways...i'm trying to do the problem without substitution because we were told to do without substitution correct?

OpenStudy (anonymous):

we can use substitution, i just wrote that after the 'u' substitution (finding a,b and c) I got stuck.. sorry if i wasnt clear!!

myininaya (myininaya):

and that whole thing is suppose to be negative not just the first integral

OpenStudy (anonymous):

I see what you did there.. ok so when I convert u back to x, the only difference is c which is 2/3

OpenStudy (anonymous):

right, i see

myininaya (myininaya):

now do you want to do sub first then the partial or the partial then do the sub?

myininaya (myininaya):

Because I did the partial first because I thought you didn't want to do the sub lol

OpenStudy (anonymous):

sub then partial.. so i can follow easier

myininaya (myininaya):

oh no that x is e^x

myininaya (myininaya):

\[-\frac{1}{3}( \int\limits\limits_{}^{}\frac{1}{e^x-1} dx +\int\limits\limits_{}^{}\frac{-e^x+1}{e^{2x}+e^x+1}dx) \]

myininaya (myininaya):

ok so if we look back at integral and not just the fraction we could do a sub and let u=e^x and so du=e^x dx which is what loser was talking about

myininaya (myininaya):

\[\int\limits_{}^{}\frac{-e^x}{e^{3x}-1} dx\]\[\int\limits_{}^{}\frac{- du}{u^3-1} =-\int\limits_{}^{}\frac{du}{(u-1)(u^2+u+1)} \]

myininaya (myininaya):

yeah.. ok so A = 1/3 B = -1/3 C= -2/3 your values in this case are correct

OpenStudy (anonymous):

first integral: (1/3)(1/(u-1))

myininaya (myininaya):

i just quoted what yu said earlier.

OpenStudy (anonymous):

awesome!

myininaya (myininaya):

\[-(\int\limits_{}^{}\frac{1}{3(u-1)} du+\int\limits_{}^{}\frac{\frac{-1}{3}u-\frac{2}{3}}{u^2+u+1} du)\]

myininaya (myininaya):

your first integral there should be the easier one to deal with...do you need help on the first one?

myininaya (myininaya):

\[-[-\frac{1}{3}\int\limits_{}^{}\frac{u}{u^2+u+1} du]-[-\frac{2}{3}\int\limits_{}^{}\frac{1}{u^2+u+1} du]\] I distribute that negative sign we have ... so on both of thse it looks like you have to do a trig sub

OpenStudy (anonymous):

no i got that one - it's -(1/3)*ln|e^x|

OpenStudy (anonymous):

wait, how did you change the first integral?

myininaya (myininaya):

what do you mean? i didn't rewrite it. i assumed you didn't need help with it. but shouldn't it be -1/3*ln|u-1|=-1/3*ln|e^x-1|

OpenStudy (anonymous):

oh nevermind, ok i and yeah it is u-1, i forgot to add that.. ok so how do you tackle the second one? do you have to do a second partial fraction?

myininaya (myininaya):

i mentioned trig sub above.

myininaya (myininaya):

can't factor that is when i hop on the trig sub bandwagon

OpenStudy (anonymous):

so you are going to complete the square and find the trig sub.. this problem is ridiculous!

OpenStudy (anonymous):

ok, let me try that first

myininaya (myininaya):

lol yep

myininaya (myininaya):

\[u^2+u+1=(u^2+u+(\frac{1}{2})^2)+1-(\frac{1}{2})^2=(u+\frac{1}{2})^2+\frac{3}{4}\]

OpenStudy (anonymous):

ok, that was killing me.. i was thinking about (u+1)^2 - u.. which wouldn't work.. but this makes sense.. and works

myininaya (myininaya):

lol if we had \[ax^2+bx+c\] on bottom you would group those parts that the variable part together \[(ax^2+bx)+c\] then factor out whatever is in front of the x^2 from both the ax^2 and b term like: \[a(x^2+\frac{b}{a}x)+c\] now the completing the square part if we add something in then we also need to take it out \[a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)+c-a(\frac{b}{2a})^2 \] so now we can write \[a(x+\frac{b}{2a})^2+c-a \frac{b^2}{4a^2}\] simplify \[a(x+\frac{b}{2a})^2+c-\frac{b^2}{4a}\]

OpenStudy (anonymous):

that will be useful.. im going to write this down.. in the meantime.. i worked on the rest and substituted (u+1/2) for h, to make it clearer so i have 1/(h^2 + 3/4)

OpenStudy (anonymous):

next to both integrals

myininaya (myininaya):

\[(u+1)^2+\frac{3}{4} =\frac{3}{4} ( \frac{4}{3} (u+1)^2+1)=\frac{3}{4}((\frac{2[u+1]}{\sqrt{3}})^2+1)\] now sub that thing inside the square

myininaya (myininaya):

by the way I'm just writing the bottom of these integrals.

myininaya (myininaya):

recall \[\tan^2(\theta)+1=\sec^2(\theta) \]

myininaya (myininaya):

so tan(theta) is the trig sub we want

OpenStudy (anonymous):

ok i'm lost here how did you get (u+1)^2+3/4=(3/4((4/3)(u+1)^2+1)

OpenStudy (anonymous):

you took 3/4 out to make the constant 1

myininaya (myininaya):

I factored out 3/4

OpenStudy (anonymous):

but why does 1/2 change in the parenthesis

myininaya (myininaya):

i wanted it 1 so i could have in the form \[\tan^2(\theta)+1=\sec^2(\theta)\]

myininaya (myininaya):

\[(u+1)^2+\frac{3}{4} \] I want something^2+1 \[\frac{4}{3} \cdot \frac{3}{4}(u+1)^2+\frac{3}{4}\] so i factored 3/4 out \[\frac{3}{4} \cdot (\frac{4}{3}(u+1)^2+1)\] that thingy inside the parenthesis is not something^2+1 \[\frac{3}{4} \cdot ([\frac{\sqrt{4}}{\sqrt{3}}(u+1)]^2+1)\]

myininaya (myininaya):

\[\frac{3}{4} ([\frac{2}{\sqrt{3}}(u+1)]^2+1)\]

myininaya (myininaya):

you can write it like this: \[\frac{3}{4}([\frac{2(u+1)}{\sqrt{3}}]^2+1)\] so it is cleary to you would you want to use as the opposite and the adjacent of some right triangle .

OpenStudy (anonymous):

you start off with (u+1/2)^2+(3/4) in the denominator.. but then i see (u+1)^2 +(3/4) which is really confusing me so the rest I can't even try to comprehend until that makes sense lol

myininaya (myininaya):

yes the 1 is suppose to be 1/2 it gets hard writing in latex sometimes is that your only question

myininaya (myininaya):

\[\frac{3}{4}([\frac{2(u+\frac{1}{2})}{\sqrt{3}}]^2+1)\] if you notice we lost that 1/2 a long time ago i have been writing 1 for the last 30 minutes on this computer thingy.

OpenStudy (anonymous):

yeah, that's why I was incredibly confused.. now it makes sense but i want to run by it again sorry

OpenStudy (anonymous):

so \[\frac{ 3 }{ 4 } (\tan^{-1}(\frac{ 2h }{ \sqrt{3} }))\]

myininaya (myininaya):

\[\text{ let } \tan(\theta)=\frac{2(u+\frac{1}{2})}{\sqrt{3}}\] \[\sqrt{3} \tan(\theta)=2u+1\] \[u=\frac{\sqrt{3} \tan(\theta)-1}{2}\] \[\sec^2(\theta) d \theta=\frac{2}{\sqrt{3}} du\]